OpenStudy (zubhanwc3):

Evaluate the integral of 2-csc^2x upper limit: pi/2 lower limit: pi/4

OpenStudy (zubhanwc3):

\[\int\limits_{\pi/4}^{\pi/2}(2-\csc^2x)dx\]